OpenStudy (anonymous):

integrate x(y^3+1)^(1/2) dydx over (x/3)<=y<=2 and 0<=x<=6

OpenStudy (anonymous):

First draw your region D.  So you want to integrate x first. To do this you need to switch the bounds. So now we have: \[0 \le x \le 3y; 0 \le y \le 2\] Giving: \[\int\limits_0^2 \int\limits_0^{3y}x(y^3+1)^{\frac{1}{2}}dx dy=\frac{1}{2}\int\limits_0^2 (3y)^2(y^3+1)^{\frac{1}{2}}dy=\frac{9}{2}\int\limits_0^1y^2(y^3+1)^{\frac{1}{2}}dy\] Now make a change of variable on y^3+1. Now it's easy.